A typist is entering text on a keyboard at a rate of 40 words per minute. If each word is 5 characters long on average, what bandwidth in bits per second between the keyboard and main memory is required to transfer the information? A space is also a character. Assume that each word is followed by one space on average.
Answer #1

Here is your answer.. Bandwith= 32 bits/sec

Here, Total word type in one minute= 40

  Each word length = 5

Each word is followed by a space so ,basically typist type 6 character (5 + 1 space) for a word

so, total character typist type in one minute = 40 * 6 = 240

[we know that size of character is 1 Byte]

[ 1 Byte = 8 bits]

So, total size of character = 240 Bytes = 240*8 bits = 1920 bits in 1 minute

bandwidth in bits per second = 1920/60 = 32 bits/sec